Sorry, you need to enable JavaScript to visit this website.

MCST Banner Image

am1weekly_2001-104_2001-110

Mission Operations Days: 2001/104 to 2001/110
April 14, 2001 00:00:00 GMT to April 20, 2001 00:00:00 GMT


Terra Spacecraft and MODIS Instrument Status:

Terra (AM-1) is in Normal Ops Mode
MODIS is in Science Mode - B Side
ComponentState(s)Comment
BlackbodyA Off; B On (290 K)Nominal
Calibration ElectronicsA Off; B OnNominal
Control ProcessorA Off; B OnNominal
Nadir DoorUnlatched, OpenNominal
Space View DoorUnlatched, OpenNominal
Solar Diffuser DoorUnlatched, ClosedNominal
FDDI FormatterA Off; B OnNominal
FDDI PortA On; B OffNominal
FIFO Memory#1 & #2 Off; #3 & #4 OnNominal
Format ProcessorA Off; B OnNominal
Power Supplies#1 Off; #2 OnNominal
PV VIS FPAA Off; B OnNominal
PV NIR FPAA Off; B OnNominal
PV SMIR FPAA Off; B OnNominal
PV LWIR FPAA Off; B OnNominal
PC LWIR FPAA Off; B OnNominal
RC Outgas HeatersAll Off (CS, IS, OS)Nominal
RC LWIR FPA HeaterOn (83 K)Nominal
RC SMIR FPA HeaterOffNominal
Scan AssemblyA Off; B OnNominal
SDSMA Off; B OffNominal
SRCAA Off; B OffNominal
PS1 Survival HeaterEnabledNominal
PS2 Survival HeaterEnabledNominal
Timing GeneratorA Off; B OnNominal
Flight SoftwareRev BD + 1 set of patches
  • FR Error Counting and Statistics (2 patches)
Nominal
Inhibit Ids SetNoneNominal
TMONs Enabled68Nominal

This Week’s Completed MODIS Activities:

Friday, April 13, 2001

2001/103 21:00:00 ATC OA-26: B-side Set Blackbody Temperature to 270K

Saturday, April 14, 2001

2001/104 15:00:00 ATC OA-26: B-side Set Blackbody Temperature to 280K
2001/104 19:00:00 ATC OA-26: B-side Set Blackbody Temperature to 285K
2001/104 22:00:00 ATC OA-26: B-side Set Blackbody Temperature to 290K

Sunday, April 15, 2001

2001/105 02:00:00 ATC OA-26: B-side Set Blackbody Temperature to 295K
2001/105 02:00:30 ATC OA-26: B-side Set Blackbody Heater duty cycle to FULL
2001/105 04:00:00 ATC OA-26: B-side Set Blackbody Temperature to 300K
2001/105 06:00:00 ATC OA-26: B-side Set Blackbody Temperature to 315K
2001/105 09:00:00 ATC OA-26: B-side Set Blackbody Temperature to 270K
2001/105 09:00:30 ATC OA-26: B-side Set Blackbody Heater duty cycle to THIRD

Monday, April 16, 2001

2001/106 05:00:00 ATC OA-26: B-side Set Blackbody Temperature to 290K

Tuesday, April 17, 2001

None

Wednesday, April 18, 2001

2001/108 21:44:03 ATC OA16: SDSM SCREENED
2001/108 23:22:55 ATC OA15: SDSM OPEN

Thursday, April 19, 2001

2001/109 01:13:28 – 01:15:59 ATC OA27: PV ECAL
2001/109 01:16:53 – 01:21:24 ATC OA28: PC ECAL

Friday, April 20, 2001

None

This Week's Scheduled MODIS Activities Not Completed:

None

Upcoming MODIS Events:

Friday, April 20, 2001

None

Saturday, April 21, 2001

None

Sunday, April 22, 2001

None

Monday, April 23, 2001

None

Tuesday, April 24, 2001

None

Wednesday, April 25, 2001

None

Thursday, April 26, 2001

None

Friday, April 27, 2001

None

In addition:

A new set of SRCA cross-talk measurements are being planned for later this month and will use memory load patches to change the size of the SRCA sector from 10 frames to 55 frames in the same manner as was done on March 2, 2001.

Maneuvers:

Again, there will be no MODIS roll maneuver in May because the roll angle would have been too great. The next attempt for a roll maneuver will be June.

There is a Drag makeup maneuver being planned tentatively for the first week in May.

MODIS Anomalies:

None.

General Instrument Comments:

MODIS is in Science Mode on the B-side with the SVD and NAD open, operating with the Formatter Resolution Flight Software Patches (FRFSP).

MODIS Telemetry Trends:

Nothing new to report.

Non-MODIS Significant Events:

None.

Limited Life Item Status:

SRCA 10W Lamp #1: 185.0 of 500 hours
SRCA 10W Lamp #2: 142.4 of 500 hours
SRCA 10W Lamp #3: 151.6 of 500 hours
SRCA 10W Lamp #4: 61.5 of 500 hours

SRCA 1W Lamp #1: 557.3 of 4000 hours
SRCA 1W Lamp #2: 276.3 of 4000 hours

Solar Diffuser Door: 1591 of 3022 Movements
Nadir Aperture Door: 532 of 1316 Movements
Space View Door: 437 of 1316 Movements